Determine the principal P that must be invested at rate
r = 3
1
2
%,
compounded monthly, so that $700,000 will be available for retirement in
t = 13
years. (Round your answer to the nearest cent.)
P = $
Answer #1
Do you really mean 312% 1nt. rate?
If this is an error, you can recalculate using the correct %. I'll
use 31.2%.
Pt = Po(1+r)^n = $700,000.

r = (31.2%/12) / 100%=0.026 = Monthly % rate expressed as a decimal.

n = 12 comp./yr * 13yrs = 156 = The #
of compounding periods.

Po(1.026)^156 = 700,000.
Po = 700000 / (1.026)^156 = $12,767.61.
answered by: macail